# ImageChain: Advancing Sequential Image-to-Text Reasoning in Multimodal Large Language Models

### v1

@@ 1 Introduction
Recent methods have extended MLLMs such as LLaVA-NeXT, Mantis (Jiang et al., 2024a), and Qwen2-VL (Wang et al., 2024b) to handle multi-image inputs. However, they typically process images independently or summarize entire scenes, rather than explicitly modeling the evolution of events over time. In contrast, sequential image reasoning requires a model to capture dependencies across frames to predict future actions. Figure 1 illustrates our proposed task, next-scene description, which consists of generating a text description of a visual scene based on a sequence of preceding frames and their corresponding descriptions – a challenge that existing models, such as LLaVA-NeXT, have yet to overcome.
@@ 5.2 Dataset and Context Length Splits
5.3 Training and Implementation Details
All of our model configurations share a common experimental backbone based on LLaVA-v1.6-Vicuna-7B (Liu et al., 2024), which aligns Vicuna (Chiang et al., 2023) and CLIP (Radford et al., 2021) into an MLLM. We fine-tune the models using low-rank adaptation (Hu et al., 2021, LoRA) in combination with DeepSpeed ZeRO-3, running approximately 1-2 hours per model on 8 NVIDIA A100 GPUs. The baseline training is conducted over 3 epochs on the StoryFrames dataset with a per-device batch size of 4 and gradient accumulation of 1. LoRA is configured with a rank $r=128$ and $\alpha=256$. A learning rate of $2\times 10^{-5}$ is applied, with a cosine learning rate schedule and a warmup ratio of 3%. For optimization, the adaptive momentum optimizer with decoupled weight decay (Loshchilov and Hutter, 2019) is used. For ImageChain trained on C2 and C3 we adjust the number of training epochs to 5, and to 7 epochs when trained on C4–7 due to the smaller training sample sizes. All other training settings remain unchanged.
@@ 5.4 Evaluation
To quantitatively assess the quality of the generated descriptions, we use similarity rate (SimRate), an adapted version of win rate (Chiang et al., 2024), as a metric. Given the inherent variability in how visual events can be described literally, conventional metrics relying on n-gram overlap are inadequate for capturing semantic equivalence (Culy and Riehemann, 2003; Bulian et al., 2022). Instead, we follow recent literature (Li et al., 2024, 2025) and adopt LLMs as evaluators (Eldan and Li, 2023). For our generated next-scene description $\hat{D}_{\tau}$, Llama 3 70B (Dubey et al., 2024) determines whether or not it conveys a similar meaning to the ground truth scene description $D_{\tau}$, effectively serving as a proxy for human evaluators (Li et al., 2023; Chiang et al., 2024). The prompt used for this framework is detailed in Appendix A. The overall SimRate follows as the fraction of comparisons where model-generated descriptions are judged semantically similar to the human-annotated ground truth descriptions.
@@
To further assess model generalization, we evaluate performance on several out-of-domain datasets that target diverse sequential reasoning challenges. Specifically, we test on three datasets derived from the Mementos benchmark (Wang et al., 2024c). Comics, which features wordless multi-panel comics; Daily-Life (DL) (Xiao et al., 2021), consisting of videos depicting everyday activities; and Robo (O’Neill et al., 2024), which contains robotics tasks. For evaluation, behavioral cues, (e.g., key verbs or verb phrases, are extracted from generated descriptions using GPT-4V (Achiam et al., 2023) and compared against human annotations using F1 score. [^2] We follow Mementos evaluation setup (Wang et al., 2024c)..
@@
MLLM 5.20 2.36 1.67 1.92 0.00 1.58 3.70
MLLM-ICL 0.19 0.34 0.00 0.00 0.00 0.00 0.20
MLLM-FT 18.11 19.59 15.83 3.85 15.38 12.13 17.52
VisualContext 12.52 16.22 11.67 1.92 15.38 8.96 13.01
FinalScene 9.63 10.47 11.67 1.92 7.69 8.44 9.71
ImageChain-NoFT 6.94 6.08 12.50 0.00 15.38 8.96 7.11
ImageChain 18.30 21.28 19.17 9.62 30.77 16.87 19.02
6 Results
@@
Table 1 shows that our proposed ImageChain model consistently achieves the highest SimRate across all evaluated context lengths, demonstrating the value of explicitly modeling a sequence of images for next-scene description task. ImageChain significantly outperforms standard fine-tuning (MLLM-FT), with a notable 15.39 percentage point improvement in SimRate on C6, the longest evaluated context length. Additionally, ImageChain achieves 19.02% on C2-6, surpassing VisualContext (13.01%) and FinalScene (9.71%), underscoring the value of incorporating both images and textual description history. Even without fine-tuning, ImageChain-NoFT outperforms MLLM on C2-6 (7.11% vs. 3.70%), demonstrating the benefits of explicitly modeling image sequences. However, fine-tuning remains essential for optimal performance, suggesting that MLLMs lack strong inherent temporal reasoning skills without additional training.
@@
We train and evaluate an ImageChain model for each context length, with results shown in Table 2. We observe that ImageChainC2-7 model outperforms the baseline ImageChainC2 model when tested on C2, achieving a 1.7 point improvement in SimRate (18.3% vs. 16.6%). This suggests that training on a broader range of context lengths helps the model generalize better, even when predicting in shorter contexts. Similarly, we find that ImageChainC2-7, which is trained on the entire range of temporal spans, consistently outperforms models trained on specific context lengths. This suggests that training across multiple context lengths enables robust sequential reasoning. Additionally, when tested on C3, ImageChainC2-7 demonstrates a substantial improvement of 5.1 points (21.3% vs. 16.2%) over the baseline ImageChainC3 model. Finally, on the C4-6 test setting, ImageChainC2-7, achieves a 2.8 point increase in SimRate (16.8% vs. 14%). This indicates that the sequential training strategy across multiple context lengths enables stronger reasoning over extended visual sequences.
Training on Longer Contexts Improves Sequential Reasoning.
Figure 5 shows the models performance for fine-tuned models trained on different context lengths. ImageChain achieves the highest overall SimRate when evaluated on C2-6, particularly when trained on longer contexts, reaching 13.6% when trained on C4-7. This surpasses other models, such as MLLM-FT, which achieves 9.8% under the same training conditions. MLLM-FT excels in short contexts but struggles with longer dependencies, suggesting limitations in handling extended sequences without explicit sequence modeling. VisualContext, underperforms on longer sequences (7.7% trained on C4-7 and evaluated on C6), highlighting the benefit of including text descriptions for fine-tuning in long contexts.
6.1 Out-of-domain Setup
Table 3 compares ImageChain with various baselines and large-scale multimodal models (GPT-4V, Gemini (Gemini Team et al., 2023)) across different domains. While GPT-4V and Gemini achieve the highest F1 scores, they operate at a much larger parameter and training data scale compared to our models, making direct comparisons challenging, but serve as a reference.
Comics.
In the comics domain, ImageChain trails Gemini by only 0.1 points (16.2 vs. 16.3) and surpasses all 7B baselines, indicating that modeling sequential dependencies benefits structured visual narratives. Figure 6 shows how ImageChain captures the evolution of events over time, unlike MLLM-FT that focus on static observations.
@@ 6.1 Out-of-domain Setup
Table 3: Zero-shot out-of-domain performance comparison (F1 score) on Comics, Daily-Life (DL), and Robotics (Robo) datasets. ImageChain excels in structured, event-based reasoning tasks (comics and robotics), trailing Gemini only by 0.1 and 0.8 points, respectively. GPT-4V and Gemini results are taken from Wang et al. (2024c).
@@
Figure 6: Comparison of next-scene descriptions. MLLM-FT (standard fine-tuning) describes static observations, while ImageChain captures sequential actions. Behavioral cues, i.e., key verbs, are highlighted.
Daily-Life.
For daily life (DL) videos, ImageChain lags behind FinalScene (20.8 vs. 26.5 F1), suggesting that the task is closer to traditional captioning since descriptions summarize videos with fewer significant changes compared to comics, where progression between frames is more explicit.
Robotics.
In robotics (Robo), ImageChain (27.1 F1) shows a substantial improvement over MLLM (15.2 F1) and MLLM-FT (14.4 F1), highlighting that explicitly modeling image sequences improves reasoning in multi-step environments, even in zero-shot out-of-domain settings. VisualContext (27.8 F1) achieves a similar performance, suggesting that leveraging prior frames without explicit text descriptions can be effective in this setting.
@@ 7 Conclusions
In this work, we introduced ImageChain, a framework designed to enhance MLLMs with explicit sequential reasoning capabilities, addressing a key limitation in existing MLLMs. Our results demonstrate substantial improvements over baseline models, with strong performance gains in both in-domain and zero-shot out-of-domain settings. Our findings highlight the importance of instruction-tuning within a multimodal, multi-turn conversation framework, suggesting promising directions for future work in refining temporal reasoning, scaling to more complex real-world scenarios and applications such as video understanding and robotics.
@@
While ImageChain enhances sequential reasoning by structuring image sequences as multi-turn conversations, it has some limitations. We test our method on LLaVA-NeXT and average the feature vectors over all frames to compute the scene-level visual representation, a simple yet effective approach. While future work can investigate performance in other models and can explore more advanced methods for combining feature vectors, this experimental setup is sufficient for evaluating our method. Scalability remains a challenge for longer sequences, as maintaining coherence across many turns may require more sophisticated techniques, which can be explored in future work. That said, our framework already models dependencies more effectively than standard MLLMs, making it a strong foundation for further improvements.
@@
This work uses publicly available licensed (CC BY 4.0) datasets consistent with their intended use (research) to ensure transparency and reproducibility. While ImageChain enhances sequential reasoning, it may inherit biases from pre-trained models and datasets. Our framework is designed for research and development purposes, and we encourage responsible use, particularly in applications involving decision-making in sensitive domains such as healthcare and robotics. We acknowledge the use of Microsoft Copilot (https://copilot.microsoft.com/) during the development of the coding experiments.
@@ Appendix A Evaluation Prompt
Figure 7: Evaluation prompt used with Llama 3 70B to annotate the predicted descriptions. The prompt structure is adapted towards description prediction from Alpaca-Eval. The positions are indicated by placeholders [$\ldots$], where the ground truth and predictions to be annotated are inserted.

### v2

@@ 1 Introduction
Recent methods have extended MLLMs such as LLaVA-NeXT, Mantis (Jiang et al., 2024a), and Qwen2-VL (Wang et al., 2024b) to handle multi-image inputs. However, they typically process images independently or summarize entire scenes, rather than explicitly modeling the evolution of events over time. In contrast, sequential image reasoning requires a model to capture dependencies across frames to predict future actions. Figure 1 illustrates our proposed task, next-scene description, which consists of generating a text description of a visual scene based on a sequence of preceding frames and their corresponding descriptions – a challenge that multimodal models, such as LLaVA-NeXT, have yet to overcome.
@@ 4 The StoryFrames Dataset
Model C2 C3 C4 C5 C6 C4–6 C2–6
MLLM 5.20 2.36 1.67 1.92 0.00 1.58 3.70
MLLM-ICL 0.19 0.34 0.00 0.00 0.00 0.00 0.20
MLLM-FT 18.11 19.59 15.83 3.85 15.38 12.13 17.52
VisualContext 12.52 16.22 11.67 1.92 15.38 8.96 13.01
FinalScene 9.63 10.47 11.67 1.92 7.69 8.44 9.71
ImageChain-NoFT 6.94 6.08 12.50 0.00 15.38 8.96 7.11
ImageChain 18.30 21.28 19.17 9.62 30.77 16.87 19.02
@@ 5.1 Models
Our primary experiments are conducted using LLaVA-v1.6 with Vicuna-7B (Chiang et al., 2023) and CLIP (Radford et al., 2021) variant (Liu et al., 2024) as the base MLLM architecture. To assess the robustness of our findings and ensure that ImageChain’s improvements generalize across backbones, we replicate our main in-domain experiments (see Section 5.2) using Qwen2-VL-7B (Wang et al., 2024b). For Qwen, we select the first frame per scene to avoid inconsistent representations from averaging, as its visual tokens are adaptively generated compared to LLaVA-1.6’s static CLIP features. Unless otherwise specified, results presented are based on the LLaVA backbone. Training and implementation details are described in Appendix C.
@@ 5.3 Evaluation
To quantitatively assess the quality of the generated descriptions, we use similarity rate (SimRate), an adapted version of win rate (Chiang et al., 2024), as a metric. Given the inherent variability in how visual events can be described literally, conventional metrics relying on n-gram overlap are inadequate for capturing semantic equivalence (Culy and Riehemann, 2003; Bulian et al., 2022). Instead, we follow recent literature (Li et al., 2024; Ziegler et al., 2024; Li et al., 2025) and adopt LLMs as evaluators (Eldan and Li, 2023). For our generated next-scene description $\hat{D}_{\tau}$, Llama 3.1 70B (Dubey et al., 2024) determines whether or not it conveys a similar meaning to the ground truth scene description $D_{\tau}$, effectively serving as a proxy for human evaluators (Li et al., 2023; Chiang et al., 2024). The prompt used for this framework is detailed in Appendix A. The overall SimRate follows as the fraction of comparisons where model-generated descriptions are judged semantically similar to the human-annotated ground truth descriptions. Section 6.1 presents a human validation study that confirms that our automated SimRate metric aligns well with human judgments.
@@
To further assess model generalization, we evaluate performance on several out-of-domain datasets that target diverse sequential reasoning challenges. Specifically, we test on three datasets derived from the Mementos benchmark (Wang et al., 2024c). Comics, which features wordless multi-panel comics; Daily-Life (DL) (Xiao et al., 2021), consisting of videos depicting everyday activities; and Robo (O’Neill et al., 2024), which contains robotics tasks. For evaluation, we follow Mementos and extract behavioral cues, (e.g., key verbs or verb phrases, from generated descriptions using GPT-4V (Achiam et al., 2023) and compared against human annotations using F1 score.
@@
MLLM 6.7 9.5 7.5 1.9 0.0 5.4 7.3
MLLM-ICL 2.3 2.7 2.5 0.0 0.0 1.6 2.3
MLLM-FT 20.0 21.3 16.7 11.5 23.1 15.7 19.6
VisualContext 13.3 17.6 16.7 9.6 38.5 16.2 15.1
FinalScene 14.1 17.6 17.5 11.5 30.8 16.8 15.6
ImageChain-NoFT 9.3 17.9 15.8 13.5 38.5 16.8 13.2
ImageChain 18.7 22.3 19.2 7.7 46.2 17.8 19.6
Table 2: SimRate (%) on StoryFrames using Qwen2-VL as an alternative backbone. ImageChain’s sequential reasoning advantages generalize across architectures. As context length increases, our explicit multi-turn formulation yields progressively stronger scores compared to standard fine-tuning (MLLM-FT).
6 Results and Analysis
6.1 In-Domain Performance
@@
Table 1 shows that our proposed ImageChain model consistently achieves the highest SimRate across all evaluated context lengths, demonstrating the value of explicitly modeling a sequence of images for next-scene description task. ImageChain also significantly outperforms standard fine-tuning (MLLM-FT), with a notable 15.39 percentage point improvement in SimRate on C6, the longest evaluated context length. Additionally, ImageChain achieves 19.02% on C2-6, surpassing VisualContext (13.01%) and FinalScene (9.71%), underscoring the value of incorporating both images and textual description history. Even without fine-tuning, ImageChain-NoFT outperforms MLLM on C2-6 (7.11% vs. 3.70%), demonstrating the benefits of explicitly modeling image sequences. However, fine-tuning remains essential for optimal performance, suggesting that MLLMs lack strong inherent temporal reasoning skills without additional training.
ImageChain Generalizes Across Backbones.
To evaluate the backbone robustness of ImageChain, we replicate our main results on Qwen2-VL-7B (Wang et al., 2024b), a multimodal model whose vision and language modules differ substantially from LLaVA-v1.6 (Liu et al., 2024).
@@
The results show a similar progression: once three or more prior scenes are provided, ImageChain dominates with +2.2% on C4-6 and +23.1% on C6 vs. standard fine-tuning, i.e., MLLM-FT. This underscores our argument that explicit multi-turn formulation is critical for extended temporal dependencies. ImageChain on Qwen likely lags behind MLLM-FT at C2 due to its instruction overhead, which offers less value in short contexts. However, as more prior scenes become available (C3+), its multi-turn conversation design increasingly benefits from modeling temporal progression. Although both methods tie overall, reflecting the dataset’s bias toward short stories (see Figure 4), ImageChain proves more effective as context length increases.
Validating SimRate With Human Alignment.
To ensure the reliability of SimRate, we conducted a human annotation study comparing Llama 3.1 70B’s judgments against human evaluations on 90 diverse examples (full description and detailed analysis in Appendix B). The Llama judge achieved 72% accuracy (95% CI: 62–80%) against human majority vote, and its overall SimRate showed no significant difference from human-aggregated ratings. While SimRate is generally aligned over all analyzed models, the Llama judge’s calibration varies by model: it was conservative for ImageChain and MLLM-FT (underestimating their human-rated scores by 10-13 pp.) and optimistic for MLLM (+7 pp.). Therefore, if ranked by human labels, ImageChain would see its score further increase by approximately 13 points.
@@
As shown in Table 3, training ImageChain on a diverse range of temporal spans (ImageChainC2-7) consistently enhances performance compared to training on fixed, specific context lengths. For instance, ImageChainC2-7 outperforms ImageChainC2 by 1.7 pp. when tested on C2 (18.3% vs. 16.6%) and surpasses ImageChainC3 by 5.1 pp. on C3 (21.3% vs. 16.2%). Even on longer contexts (C4-6), ImageChainC2-7 shows a 2.8 point gain over ImageChainC4-7 (16.8% vs. 14.0%). These results highlight that exposure to varied temporal spans during training enables more robust sequential reasoning and better generalization across different context lengths. Further analyses in Appendix D show that when models are trained on longer contexts (C4-7), ImageChain demonstrates superior or highly competitive performance against baselines on both short and long evaluation contexts.
@@
Table 4: Out-of-domain performance (F1 score) on Comics, Daily-Life (DL), and Robotics (Robo) datasets. ImageChain excels in structured, event-based reasoning tasks such as comics, trailing Gemini only by 0.1. : results are taken from Wang et al. (2024c).
6.2 Out-of-Domain Performance
Table 4 shows ImageChain’s out-of-domain performance against various MLLMs and video models. Larger API-based models (GPT-4V, Gemini (Gemini Team et al., 2023)) are included as references, though their scale makes direct comparison challenging.
Comics. In the comics domain, ImageChain trails Gemini by only 0.1 points (16.2 vs. 16.3) and surpasses all 7B baselines including video models, indicating that modeling sequential dependencies benefits structured visual narratives. Figure 5 shows how ImageChain captures the evolution of events over time, unlike MLLM-FT that focus on static observations.
@@ 6.2 Out-of-Domain Performance
Figure 5: Comparison of descriptions for sequential images in an out-of-domain setting (comics). While MLLM-FT (standard fine-tuning) tends to describe static observations, ImageChain captures dynamic, sequential actions. Key behavioral cues, such as action-oriented verbs, are highlighted in the text.
Daily-Life. For daily life (DL) videos, ImageChain lags behind FinalScene (20.8 vs. 26.5 F1), suggesting that the task is closer to traditional captioning since descriptions summarize videos with fewer significant changes compared to comics, where progression between frames is more explicit.
Robotics. In robotics (Robo), ImageChain (27.1 F1) shows a substantial improvement over MLLM (15.2 F1) and MLLM-FT (14.4 F1) as well as both video models (20.1 and 24.9 F1), highlighting that explicitly modeling image sequences improves reasoning in multi-step environments, even in out-of-domain settings. VisualContext (27.8 F1) achieves a similar performance, suggesting that leveraging prior frames without explicit text descriptions can be effective in this setting.
@@ 7 Conclusions
In this work, we introduced ImageChain, a framework designed to enhance MLLMs with explicit sequential reasoning capabilities, addressing a key limitation in existing MLLMs. Our results demonstrate substantial improvements over baseline models, with strong performance gains in both in-domain and zero-shot out-of-domain settings. Our findings highlight the importance of instruction-tuning within a multimodal, multi-turn conversational framework. This suggests promising directions for future work in refining temporal reasoning, scaling to more complex scenarios, and applications like video understanding or assistive robotics.
@@
While ImageChain enhances sequential reasoning by structuring image sequences as multi-turn conversations, maintaining coherence across many turns may require more sophisticated techniques, which can be explored in future work. That said, our framework already models dependencies more effectively than standard MLLMs, making it a strong foundation for further improvements.
@@
This work uses publicly available licensed (CC BY 4.0) datasets consistent with their intended use (research) to ensure transparency and reproducibility. While ImageChain enhances sequential reasoning, it may inherit biases from pre-trained models and datasets. Our framework is designed for research and development purposes, and we encourage responsible use, particularly in applications involving decision-making in sensitive domains such as healthcare and robotics. We acknowledge the use of Microsoft Copilot (https://copilot.microsoft.com/) during the development of the coding experiments. To ensure ethical data annotation, we recruited participants, balancing gender. Annotators provided informed consent and were compensated fairly at a rate of £9 per hour. Each participant annotated 12 examples (see Appendix B) using a task interface designed for clarity and ease of use. The task involved no sensitive or harmful content, and all data used were synthetic or publicly available, with no personally identifiable information involved.
@@ Appendix A Evaluation Prompt
Figure 6: Evaluation prompt used with Llama 3.1 70B to annotate the predicted descriptions. The prompt structure is adapted towards description prediction from Alpaca-Eval. The positions are indicated by placeholders [$\ldots$], where the ground truth and predictions to be annotated are inserted.
Appendix B Human Annotation Study
Study Design.
Ninety examples were selected for evaluation, stratified by context length and sampled proportionally from the outputs of ImageChain, MLLM, and MLLM-FT. Three human annotators were recruited via Prolific (https://www.prolific.com/), ensuring a balanced gender distribution and a minimum education level of an undergraduate degree. Annotators provided informed consent and were compensated at a rate of £9 per hour, in line with fair pay guidelines. Each participant annotated 12 examples, which included three gold-standard control examples (labeled by the authors for quality control; annotators who failed these controls were excluded from the final analysis). For each example, annotators compared the ground truth next-scene description with a model-generated description, rating the semantic overlap on a five-point Likert scale (1 = “completely different meanings”, 5 = “essentially identical meanings”). The instructions and the annotation interface are shown in Figure 7 and Figure 8 respectively.
Figure 7: Instructions of human annotation study
Figure 8: Example of the annotation interface
Data Analysis and Inter-Annotator Agreement.
For subsequent quantitative analysis (Hallgren, 2012), the 5-point Likert scores were binarized: scores $\geq$ 3 were considered as positive class. This threshold resulted in a positive class prevalence of 23% in the human-annotated dataset. The reference label for each example was determined by a majority vote among the three annotators.
Inter-annotator agreement is fair (Landis and Koch, 1977), with Fleiss’ $\kappa$ = 0.30 (Fleiss, 1971) and Krippendorff’s $\alpha_{\text{ordinal}}=0.32$ (Krippendorff, 2018) when computed on the original 1-to-5 ranks, indicating moderate consistency in how annotators judge semantic similarity, even if they differ on the exact rating level.
Alignment with Llama Judge.
Compared to human annotations, the Llama 3.1 70B judge achieves 72% accuracy, Wilson (Wilson, 1927) 95% CI: 62–80%, with balanced errors (10 false positives, 15 false negatives). McNemar’s exact test for directional bias (McNemar, 1947) produced a p-value of 0.42, indicating no significant systematic bias in one direction over the other. Overall, the SimRate produced by the Llama 3.1 70B judge differed from the human-aggregated SimRate by 5.6 percentage points (95% CI for the difference: –17.3 pp to +6.2 pp). Since this confidence interval includes zero, there is no statistically significant difference between the overall SimRate produced by the Llama judge and human evaluators.
Model-Specific Calibration Differences.
While the overall alignment was accurate, the Llama 3.1 70B judge exhibited some model-specific calibration differences when compared to human ratings:
  • For ImageChain outputs: The Llama judge was conservative. It achieved 87% accuracy on these samples but tended to underestimate positive cases. The human-rated SimRate for ImageChain was 13 percentage points higher than the Llama-judged SimRate.
  • For MLLM outputs: The Llama judge was slightly liberal. It achieved 67% accuracy and tended to overestimate positive cases. The human-rated SimRate for MLLM was 7 percentage points lower.
  • For MLLM-FT outputs: The Llama judge was again conservative. It achieved 67% accuracy, and the human-rated SimRate for MLLM-FT was 10 percentage points higher.
These differences suggest that if the main results were ranked purely based on human labels, ImageChain and MLLM-FT would likely see their scores increase relative to the Llama-judged SimRates, while MLLM’s score would decrease.
Appendix C Training and Implementation Details
We fine-tune both LLaVA-v1.6 (Liu et al., 2024) and Qwen2-VL-7B (Wang et al., 2024b) using low-rank adaptation (Hu et al., 2021, LoRA) in combination with DeepSpeed ZeRO-3, running approximately 1-2 hours per model on 8 NVIDIA A100 GPUs. The baseline training is conducted over 3 epochs on the StoryFrames dataset with a per-device batch size of 4 and gradient accumulation of 1. LoRA is configured with a rank $r=128$ and $\alpha=256$. A learning rate of $2\times 10^{-5}$ is applied, with a cosine learning rate schedule and a warmup ratio of 3%. For optimization, the adaptive momentum optimizer with decoupled weight decay (Loshchilov and Hutter, 2019) is used. For ImageChain trained on C2 and C3 we adjust the number of training epochs to 5, and to 7 epochs when trained on C4–7 due to the smaller training sample sizes. All other training settings remain unchanged.
Appendix D Context Length Ablation
Figure 9 shows the models performance for fine-tuned models trained on different context lengths. ImageChain achieves the highest overall SimRate when evaluated on C2-6, particularly when trained on longer contexts, reaching 13.6% when trained on C4-7. This surpasses other models, such as MLLM-FT, which achieves 9.8% under the same training conditions. MLLM-FT excels in short contexts but struggles with longer dependencies, suggesting limitations in handling extended sequences without explicit sequence modeling. VisualContext under-performs on longer sequences (7.7% trained on C4-7 and evaluated on C6), highlighting the benefit of including text descriptions for fine-tuning in long contexts.